OpenStudy (anonymous):

Can someone tell me exactly why there can't be two y-values for one x-value in a function?

OpenStudy (anonymous):

Because that one guy who wrote that one book said so... lol just kidding.. I have no idea

OpenStudy (anonymous):

lol thanks

jimthompson5910 (jim_thompson5910):

Think of the simple function f(x,y) = x+y Note: it's perfectly possible to have two inputs This is a fancy version of the addition operation we're all used to. So for instance, f(1, 2) = 1+2 = 3 But if f(1,2) = some other number, then the plus operator is not well defined In other words, we cannot predict what the result of 1+2 is if more than one output occurs So that is why one output is very important

OpenStudy (anonymous):

can you explain it like in a graph or something? no? because I didn't quite understand. I mean you do know what you're talking about hehe

OpenStudy (anonymous):

look at the power point you will see it clearly explained